### v1

@@
22email: antonino.nunnari@inaf.it 33institutetext: Department of Astronomy & McDonald Observatory, The University of Texas at Austin, 2515 Speedway, Austin, TX 78712, USA 44institutetext: ASI - Space Science Data Center, via del Politecnico snc, I-00133 Roma, Italy 55institutetext: Materials Science and Applied Mathematics, Malmö University, SE-205 06 Malmö, Sweden 66institutetext: Institute for Astronomy, University of Hawai’i at Manoa, Honolulu, HI 96822, USA 77institutetext: LMU München, Universitätssternwarte, Scheinerstr. 1, 81679 München, Germany 88institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 99institutetext: INAF – Osservatorio Astrofisico di Torino, via Osservatorio 20,10025 Pino Torinese (TO), Italy 1010institutetext: Department of Physics and Astronomy, Texas Christian University, TCU Box 298840 Fort Worth, TX 76129, USA 1111institutetext: INAF – Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, via Piero Gobetti 93/3, 40129 Bologna, Italy 1212institutetext: Dipartimento di Fisica, Sapienza Università di Roma, P.le A. Moro 5, I-00185 Roma, Italy 1313institutetext: INAF - Osservatorio Astronomico di Capodimonte, salita Moiariello 16, 80131, Naples, Italy 1414institutetext: Department of Science and Technology, Parthenope University of Naples, CDN-IC4, Naples, I-80143, Italy 1515institutetext: Dipartimento di Fisica e Astronomia, Università degli Studi di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1616institutetext: Department of Astronomy, School of Science, The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan 1717institutetext: IAC- Instituto de Astrofísica de Canarias, Calle Vía Lactea s/n, E-38205 La Laguna, Tenerife, Spain 1818institutetext: Departmento de Astrofísica, Universidad de La Laguna, E-38206 La Laguna, Tenerife, Spain 1919institutetext: Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Laboratoire Lagrange, France
Classical Cepheids in the Galactic thin disk.
@@
Classical Cepheids (CCs) have long been considered excellent tracers of the chemical evolution of the Milky Way’s young disk. We present a homogeneous, Non-Local Thermodynamical Equilibrium (NLTE) spectroscopic analysis of 401 Galactic CCs, based on 1,351 high-resolution optical spectra, spanning Galactocentric distances from 4.6 to 29.3 kpc. Using PySME with MARCS atmospheres and state-of-the-art grids of NLTE departure coefficients, we derive atmospheric parameters and abundances for key species tracing multiple nucleosynthetic channels (O, Na, Mg, Al, Si, S, Ca, Ti, Mn, Fe, Cu). Our sample–the largest CC NLTE dataset to date– achieves high internal precision and enables robust modeling of present-day thin-disk abundance patterns and radial gradients.
@@
We estimate abundance gradients using three analytic prescriptions (linear, logarithmic, bilinear with a break) within a Bayesian, outlier-robust framework, and we also apply Gaussian Process Regression to capture non-parametric variations. We find that NLTE atmospheric parameters differ systematically from LTE determinations. Moreover, iron and most elemental abundance profiles are better described by non-linear behavior rather than by single-slope linear models: logarithmic fits generally outperform simple linear models, while bilinear fits yield inconsistent break radii across elements. Gaussian Process models reveal a consistent outer-disk flattening of [X/H] for nearly all studied elements. The [X/Fe] ratios are largely flat with Galactocentric radius, indicating coherent chemical scaling with iron across the thin disk, with modest positive offsets for Na and Al and mild declines for Mn and Cu.
@@
Finally, Cepheid kinematics confirm thin-disk orbits for the great majority of the sample. Comparison with recent literature shows overall agreement but highlights NLTE-driven differences, especially in outer-disk abundances. These results provide tighter empirical constraints for chemo-dynamical models of the Milky Way and set the stage for future NLTE mapping with upcoming large spectroscopic surveys.
@@
Classical Cepheids – Galaxy formation and evolution – Spectroscopy – Chemical Abundances – …1 Introduction
@@
Radial abundance gradients play a key role in constraining Galactic chemo-dynamical models (palla20; spitoni23; lemasle22). The radial trends observed in the present-day chemical abundances of the thin disk, when compared with theoretical predictions across different Galactocentric distances and stellar ages, allow us to constrain its chemical enrichment history (chiappini01). Furthermore, radial gradients and their steady flattening as a function of time soundly support stellar radial migrations in shaping the current metallicity distribution function of the thin disk (hou00; prantzos23). Moreover, they also provide an independent constraint on the star formation efficiency when moving from the innermost to the outermost disk regions (spitoni23).
@@
Radial gradients across the thin disk have been investigated by using different stellar tracers: open clusters (friel02; magrini23; myers22; occaso24; otto25; dalponte25), planetary nebulae (Maciel13; stanghellini18), HII regions (fernandez17), and OB stars (daflon04; nieva12; braganca19). Classical Cepheids (CCs) have also been extensively used to investigate metallicity distributions and chemical gradients in the Galactic thin disk (yong06; szildi07; romaniello2008influence; pedicelli2010new; luck11; lucklambert11; lemasle2013galactic, lemasle2017detailed; genovali2013metallicity, genovali2014fine, genovali15; korotin14; martin15; andrievsky16; da2016neutron, da2022new, da2023oxygen; proxauf2018new; luck18; inno2019first; kovtyukh22; ripepi22; trentin23, Trentin24B, trentin24A, trentin25). Cepheids are radially pulsating variable stars that serve as excellent distance indicators, since their individual distances can be estimated with an accuracy of 1-3%. They also trace young stellar populations (bono2024cepheids). CCs are central helium-burning, intermediate-mass stars younger than a few hundred million years, with pulsation periods ranging from days to roughly one year. They are ubiquitous across the thin disk and the Galactic center (matsunaga11; bono2024cepheids).
@@
Since CCs are typically distributed across the thin disk, their intrinsic colors are affected by uncertainties due to reddening corrections. This is the main reason why atmospheric parameters are estimated directly from the spectra. Several approaches have been proposed to address this issue. In particular, proxauf2018new and da2022new adopted the line depth ratio method (kovtyukh07) based on empirical calibrations of different element pairs to derive effective temperature (kovtyukh07; elgueta24). A new temperature scale for Galactic CCs based on a data-driven, machine-learning technique applied to observed spectra, was recently suggested by lemasle20 in which the flux ratios of different spectral features were tied to the effective temperatures derived using the infrared surface-brightness method (hanke18).
@@
A significant fraction of abundance analysis of CCs is based on Local Thermodynamic Equilibrium (LTE) atmosphere models, such as those of castelli2004new and gustafsson2008grid. However, both theoretical predictions and empirical evidence show that LTE modeling can be prone to possible systematics (lind24; bergemann25). These systematics become more relevant when moving from dwarf stars to giants, and from metal-rich to metal-poor populations (e.g., thevenin1999stellar; idiart2000non; bergemann2012non; hansen2013lte; fabrizio2021use). Moreover, NLTE corrections are element- and line-dependent, with neutral lines more affected than ionized lines (kiselman2001nlte; collet2005effects; merle2011grid; tautvaivsiene2015gaia; duffau2017gaia).
@@
Chemical abundances for CCs based on a NLTE approach have been provided by andrievsky16 for all their analyzed elements, by luck11 for CNO elements, by martin15 for Fe and O, and by korotin14 for O. However, in these investigations the stellar atmospheric parameters were derived by using an LTE approach.
@@
Recently, amarsi2020galah provided NLTE departure coefficients for H, Li, C, N, O, Na, Mg, Al, Si, K, Ca, Mn and Ba. Additionally, NLTE departure coefficients are provided by mallinson24 for Ti, by amarsi22 for Fe, by caliskan25 for Cu and by amarsi25 for S. These departure coefficient are computed on a grid of 3756 1D MARCS model atmospheres (gustafsson2008grid) covering , , and dex. These comprehensive grids allow proper spectral modeling and NLTE analysis.
@@
In this study, we apply a complete NLTE approach—deriving both atmospheric parameters and chemical abundances—for Galactic CCs by using high–resolution optical spectra. Although the statistical sampling in the inner (Galactocentric distance, ) and in the outer () disk is sparse, there is solid empirical evidence for radial gradients for most of the elements that have been investigated. More specifically, our dataset spans a wide range in Galactocentric distances (=5–29 kpc) and provides homogeneous NLTE abundances for light (O), odd-Z (Na, Al, Cu), (Mg, S, Si, Ca, Ti), and iron peak (Mn, Fe) elements. The current analysis is a stepping stone for exploiting thousands of new high-resolution spectra for CCs that will be collected from upcoming large spectroscopic surveys in optical (dejong16, 4MOST) and near-infrared (cirasuolo12, MOONS). This paper is organized as follows: in § 2, we present the spectral dataset and the line list we adopted for the spectral analysis. In § 3 we show the results concerning the atmospheric parameters and the chemical abundances, including the gradients, and the kinematic properties. The summary of the results, the discussion and a brief outline of the near future developments of this project are presented in § 4. The paper also includes several appendices. Appendix A presents the tables included in the paper, while Appendix B discusses the new distances of two CCs. Appendix C provides details of the adopted spectral synthesis, and Appendix D presents the validation of NLTE atmospheric parameters and chemical abundances.
@@ 2 Dataset and method
The current spectroscopic sample includes 401 CCs distributed across the thin disk, as shown in Fig. 1111NASA/JPL-Caltech/R. Hurt (SSC/Caltech), https://milkyway-plot.readthedocs.io/en/. Among them, 379 CCs have already been collected and discussed in da2023oxygen (hereinafter da2023oxygen). We added 66 High-Resolution (HR) spectra for 22 CCs (proprietary plus public archives): 10 with and 4 with .
@@ 2.1 Spectral sample
In the spectroscopic sample presented in da2023oxygen, individual heliocentric distances were estimated using the following methods: Gaia DR3 parallaxes (72%), W1-band Period-Luminosity (PL) relations (25%), K-band PL (2%), and J-band PL (1%). This ranking was adopted in da2023oxygen to favor purely geometrical distances over distances from PL relations. On the other hand, the ranking of PLs is W1-K-J to minimize the impact of extinction, which is less severe at redder passbands. The da2023oxygen sample includes 1,285 HR spectra and more than 80% have a signal-to-noise ratio per pixel greater than 100. For a comprehensive description of the da2023oxygen spectroscopic sample the reader is referred to da2022new and da2023oxygen.
The da2023oxygen dataset was complemented with 22 CCs for which 40 HARPS-N, 10 UVES and 16 ESPaDOnS spectra were available from public archives: 35 HARPS-N are proprietary spectra collected in the Stellar Population Astrophysics222The Stellar Population Astrophysics is an ongoing project based on a Large Program conducted at the Telescopio Nazionale Galileo (TNG) using the HARPS-N and GIANO B echelle spectrographs for about 74 nights from 2018 to 2021 (origlia19) program, while the other 5 were available in the TNG database and analyzed in ripepi21. The additional UVES and ESPaDOnS spectra have been analyzed in trentin23; Trentin24B, martin15, andrievsky16 and kovtyukh22. Details of the spectroscopic sample, such as the spectrographs, the resolution, and the spectral coverage, are provided in Table 1. We estimated the individual distances by using the same approach adopted by da2023oxygen: Gaia EDR3 parallaxes for 20 CCs and W1-band PL for 2 CCs. For two CCs (ASAS 181024–2049.6 and ASAS SN J065046.50–085808.7) we estimated a different distance from andrievsky16 and trentin23. The new distance estimates are discussed in more detail in the Appendix B.
@@ 2.2 Spectral synthesis method
We utilized a Python version of Spectroscopy Made Easy (SME, piskunov17) for spectral synthesis, i.e. PySME333pysme-astro=0.6.20, https://github.com/MingjieJian/SME (wehrhahn21). PySME generates synthetic spectra based on a given set of atmospheric parameters, specified spectral intervals, and spectral resolution. It determines the optimal values for the selected atmospheric parameters by fitting the synthetic spectra to the observed data, taking into account the data uncertainties. The free parameters in the fitting process can include one or more stellar parameters, specific elemental abundances, and parameters related to atomic transitions in the line list. PySME provides two uncertainties according to two different methods. The first is based on SME statistics, i.e. uses a metric based on the distribution of the derivatives for each free parameter, estimating the cumulative distribution function of the generalized normal distribution (piskunov17). The second is based solely on the least-squares fit. Further details on PySME are presented in Appendix C.
@@
Fe I, Fe II, Ti I and Ti II lines. We first tested this line list in a NARVAL

### v2

@@
22email: antonino.nunnari@inaf.it 33institutetext: Department of Astronomy & McDonald Observatory, The University of Texas at Austin, 2515 Speedway, Austin, TX 78712, USA 44institutetext: ASI - Space Science Data Center, via del Politecnico snc, I-00133 Roma, Italy 55institutetext: Materials Science and Applied Mathematics, Malmö University, SE-205 06 Malmö, Sweden 66institutetext: Institute for Astronomy, University of Hawai’i at Manoa, Honolulu, HI 96822, USA 77institutetext: LMU München, Universitätssternwarte, Scheinerstr. 1, 81679 München, Germany 88institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 99institutetext: INAF – Osservatorio Astrofisico di Torino, via Osservatorio 20,10025 Pino Torinese (TO), Italy 1010institutetext: Department of Physics and Astronomy, Texas Christian University, TCU Box 298840 Fort Worth, TX 76129, USA 1111institutetext: INAF – Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, via Piero Gobetti 93/3, 40129 Bologna, Italy 1212institutetext: Dipartimento di Fisica, Sapienza Università di Roma, P.le A. Moro 5, I-00185 Roma, Italy 1313institutetext: INAF - Osservatorio Astronomico di Capodimonte, salita Moiariello 16, 80131, Naples, Italy 1414institutetext: Department of Science and Technology, Parthenope University of Naples, CDN-IC4, Naples, I-80143, Italy 1515institutetext: Dipartimento di Fisica e Astronomia, Università degli Studi di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1616institutetext: Department of Astronomy, School of Science, The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan 1717institutetext: IAC- Instituto de Astrofísica de Canarias, Calle Vía Lactea s/n, E-38205 La Laguna, Tenerife, Spain 1818institutetext: Departmento de Astrofísica, Universidad de La Laguna, E-38206 La Laguna, Tenerife, Spain 1919institutetext: Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Laboratoire Lagrange, France 2020institutetext: Astronomical Observatory, Odessa National University, Shevchenko Park, 65014, Odessa, Ukraine
Classical Cepheids in the Galactic thin disk
@@
Classical Cepheids (CCs) have long been considered excellent tracers of the chemical evolution of the Milky Way’s young disk. We present a homogeneous, non-local thermodynamical equilibrium (NLTE) spectroscopic analysis of 401 Galactic CCs, based on 1,351 high-resolution optical spectra, spanning Galactocentric distances from 4.6 to 29.3 kpc. Using PySME with MARCS atmospheres and state-of-the-art grids of NLTE departure coefficients, we derived the atmospheric parameters and abundances for key species tracing multiple nucleosynthetic channels (O, Na, Mg, Al, Si, S, Ca, Ti, Mn, Fe, and Cu). Our sample is the largest CC NLTE dataset to date and it achieves high internal precision, enabling the robust modelling of present-day thin-disk abundance patterns and radial gradients.
@@
We estimate abundance gradients using three analytic prescriptions (linear, logarithmic, bilinear with a break) within a Bayesian, outlier-robust framework. We also applied Gaussian process (GP) regression to capture non-parametric variations. We find that NLTE atmospheric parameters differ systematically from LTE determinations. Moreover, iron and most elemental abundance profiles are better described by non-linear behaviour rather than by single-slope linear models: logarithmic fits generally outperform simple linear models, while bilinear fits yield inconsistent break radii across elements. GP models reveal a consistent outer-disk flattening of [X/H] for nearly all studied elements. The [X/Fe] ratios are largely flat with Galactocentric radius, indicating coherent chemical scaling with iron across the thin disk, with modest positive offsets for Na and Al and mild declines for Mn and Cu.
@@
Finally, Cepheid kinematics confirm thin-disk orbits for the great majority of the sample. Comparisons with recent literature shows an overall agreement, while also highlighting NLTE-driven differences, especially in outer-disk abundances. These results provide tighter empirical constraints for chemo-dynamical models of the Milky Way and set the stage for future NLTE mapping with upcoming large spectroscopic surveys.
@@
Stars: variables: Cepheids – Galaxy: abundances – Galaxy: disk – Techniques: spectroscopic1 Introduction
@@
Radial abundance gradients play a key role in constraining Galactic chemo-dynamical models (Palla et al. 2020; Spitoni et al. 2023; Lemasle et al. 2022). The radial trends observed in the present-day chemical abundances of the thin disk, when compared with theoretical predictions across different Galactocentric distances and stellar ages, allow us to constrain its chemical enrichment history (Chiappini et al. 2001). Furthermore, radial gradients and their steady flattening as a function of time soundly support stellar radial migrations in shaping the current metallicity distribution function of the thin disk (Hou et al. 2000; Prantzos et al. 2023). Moreover, they also provide an independent constraint on the star formation efficiency when moving from the innermost to the outermost disk regions (Spitoni et al. 2023).
@@
Radial gradients across the thin disk have been investigated by using different stellar tracers: open clusters (Friel et al. 2002; Magrini et al. 2023; Myers et al. 2022; Carbajo-Hijarrubia et al. 2024; Otto et al. 2025; Dal Ponte et al. 2025), planetary nebulae (Maciel and Costa 2013; Stanghellini and Haywood 2018), HII regions (Fernández-Martín et al. 2017), and OB stars (Daflon and Cunha 2004; Nieva and Przybilla 2012; Bragança et al. 2019). Classical Cepheids (CCs) have also been extensively used to investigate metallicity distributions and chemical gradients in the Galactic thin disk (Yong et al. 2006; Sziládi et al. 2007; Romaniello et al. 2008; Pedicelli et al. 2010; Luck et al. 2011; Luck and Lambert 2011; Lemasle et al. 2013, 2017; Genovali et al. 2013, 2014, 2015; Korotin et al. 2014; Martin et al. 2015; Andrievsky et al. 2016; da Silva et al. 2016, 2022, 2023; Proxauf et al. 2018; Luck 2018; Inno et al. 2019; Kovtyukh et al. 2022; Ripepi et al. 2022; Trentin et al. 2023, 2024a, 2024b, 2025). Cepheids are radially pulsating variable stars that serve as excellent distance indicators, since their individual distances can be estimated with an accuracy of 1-3%. They also trace young stellar populations (Bono et al. 2024). CCs are central helium-burning, intermediate-mass stars younger than a few hundred million years, with pulsation periods ranging from days to roughly one year. They are ubiquitous across the thin disk and the Galactic centre (Matsunaga et al. 2011; Bono et al. 2024).
@@
Since CCs are typically distributed across the thin disk, their intrinsic colours are affected by uncertainties due to reddening corrections. This is the main reason why atmospheric parameters are estimated directly from the spectra. Several approaches have been proposed to address this issue. In particular, Proxauf et al. (2018) and da Silva et al. (2022) adopted the line depth ratio method (Kovtyukh 2007) based on empirical calibrations of different element pairs to derive effective temperature (Kovtyukh 2007; Elgueta et al. 2024). A new temperature scale for Galactic CCs based on a data-driven, machine-learning technique applied to observed spectra was recently suggested by Lemasle et al. (2020). Specifically, the flux ratios of different spectral features are tied to the effective temperatures derived using the infrared surface-brightness method (Hanke et al. 2018).
@@
A significant fraction of abundance analysis of CCs is based on local thermodynamic equilibrium (LTE) atmosphere models, such as those of Castelli and Kurucz (2004, ATLAS) and Gustafsson et al. (2008, MARCS). However, both theoretical predictions and empirical evidence show that LTE modelling can be prone to possible systematics (Lind and Amarsi 2024; Bergemann and Hoppe 2025). These systematics become more relevant when moving from dwarf stars to giants, and from metal-rich to metal-poor populations (e.g., Thévenin and Idiart 1999; Idiart and Thévenin 2000; Bergemann et al. 2012; Hansen et al. 2013; Fabrizio et al. 2021). Moreover, NLTE corrections are element- and line-dependent, with lines of neutral atoms being more affected than those of ionized species (Kiselman 2001; Collet et al. 2005; Merle et al. 2011; Tautvaišienė et al. 2015; Duffau et al. 2017).
@@
Chemical abundances for CCs based on a NLTE approach have been provided by Andrievsky et al. (2016) for all their analysed elements, as well as by Luck et al. (2011) for CNO elements, by Martin et al. (2015) for O, S and Ba, and by Korotin et al. (2014) for O. However, in these investigations the stellar atmospheric parameters were derived by using an LTE approach.
@@
Recently, Amarsi et al. (2020) provided NLTE departure coefficients for H, Li, C, N, O, Na, Mg, Al, Si, K, Ca, Mn, and Ba. Additionally, NLTE departure coefficients are provided by Mallinson et al. (2024) for Ti, by Amarsi et al. (2022) for Fe, by Caliskan et al. (2025) for Cu and by Amarsi et al. (2025) for S. These departure coefficient are computed on a grid of 3756 1D MARCS model atmospheres (Gustafsson et al. 2008) covering , , and dex. These comprehensive grids allow proper spectral modelling and NLTE analysis.
@@
In this study, we apply a complete NLTE approach–deriving both atmospheric parameters and chemical abundances–for Galactic CCs by using high-resolution optical spectra. Although the statistical sampling in the inner (Galactocentric distance, ) and in the outer () disk is sparse, there is solid empirical evidence of radial gradients for most of the elements that have been investigated. More specifically, our dataset spans a wide range in Galactocentric distances (=5–29 kpc) and provides homogeneous NLTE abundances for light (O), odd-Z (Na, Al, Cu), (Mg, Si, S, Ca, Ti), and iron peak (Mn, Fe) elements. The current analysis is a stepping stone for exploiting thousands of new high-resolution spectra for CCs that will be collected from upcoming large spectroscopic surveys in optical (de Jong et al. 2016, 4MOST) and near-infrared (Cirasuolo et al. 2012, MOONS). This paper is organized as follows: in Sect. 2, we present the spectral dataset and the line list we adopted for the spectral analysis. In Sect. 3 we show the results concerning the atmospheric parameters and the chemical abundances, including the gradients, and the kinematic properties. The summary of the results, the discussion and a brief outline of the near future developments of this project are presented in Sect. 4. The paper also includes several appendices. Appendix A presents the tables included in the paper, while Appendix B discusses the new distances of two CCs. Appendix C provides details of the adopted spectral synthesis, and Appendix D presents the validation of NLTE atmospheric parameters and chemical abundances.
@@ 2 Dataset and method
The current spectroscopic sample includes 401 CCs distributed across the thin disk, as shown in Fig. 1111NASA/JPL-Caltech/R. Hurt (SSC/Caltech), https://milkyway-plot.readthedocs.io/en/. Among them, 379 CCs have already been collected and discussed in da Silva et al. (2023) (hereinafter dS23). We added 66 High-Resolution (HR) spectra for 22 CCs (proprietary plus public archives): 10 with and 4 with .
@@ 2.1 Spectral sample
In the spectroscopic sample presented in dS23, individual heliocentric distances were estimated using the following methods: Gaia DR3 parallaxes prior-weighed according to Bailer-Jones et al. (2021), which also includes the Lindegren et al. (2021) bias (72%), W1-band Period-Luminosity (PL) relations (25%), K-band PL (2%), and J-band PL (1%). This ranking was adopted in dS23 to favor purely geometrical distances over distances from PL relations. On the other hand, the ranking of PLs is W1-K-J to minimize the impact of extinction, which is less severe at redder passbands. The dS23 sample includes 1,285 HR spectra and more than 80% have a signal-to-noise ratio per pixel greater than 100. For a comprehensive description of the dS23 spectroscopic sample the reader is referred to da Silva et al. (2022) and dS23.
The dS23 dataset was complemented with 22 CCs for which 40 HARPS-N, 10 UVES and 16 ESPaDOnS spectra were available from public archives: 35 HARPS-N are proprietary spectra collected in the Stellar Population Astrophysics222The Stellar Population Astrophysics is an ongoing project based on a Large Programme conducted at the Telescopio Nazionale Galileo (TNG) using the HARPS-N and GIANO B echelle spectrographs for about 74 nights from 2018 to 2021 (Origlia et al. 2019) programme, while the other 5 were available in the TNG database and analysed in Ripepi et al. (2021). The additional UVES and ESPaDOnS spectra have been analysed in Trentin et al. (2023, 2024a), Martin et al. (2015), Andrievsky et al. (2016) and Kovtyukh et al. (2022). Details of the spectroscopic sample, such as the spectrographs, the resolution, and the spectral coverage, are provided in Table 1. We estimated the individual distances by using the same approach adopted by dS23: Gaia EDR3 parallaxes for 20 CCs and W1-band PL for 2 CCs. For two CCs (ASAS 181024-2049.6 and ASAS SN J065046.50-085808.7) we estimated a different distance from Andrievsky et al. (2016) and Trentin et al. (2023). The new distance estimates are discussed in more detail in the Appendix B.
@@ 2.2 Spectral synthesis method
We utilized a Python version of Spectroscopy Made Easy (SME, Piskunov and Valenti 2017) for spectral synthesis, i.e. PySME333pysme-astro=0.6.20, https://github.com/MingjieJian/SME (Wehrhahn 2021). PySME generates synthetic spectra based on a given set of atmospheric parameters, specified spectral intervals, and spectral resolution. It determines the optimal values for the selected atmospheric parameters by fitting the synthetic spectra to the observed data, taking into account the data uncertainties. The free parameters in the fitting process can include one or more stellar parameters, specific elemental abundances, and parameters related to atomic transitions in the line list. PySME provides two uncertainties according to two different methods. The first is based on SME statistics, i.e. uses a metric based on the distribution of the derivatives for each free parameter, estimating the cumulative distribution function of the generalized normal distribution (Piskunov and Valenti 2017). The second is based solely on the least-squares fit. Further details on PySME are presented in Appendix C.
@@
Fe I, Fe II, Ti I and Ti II lines. Since CCs are intrinsically variable stars, they do not have fixed benchmark atmospheric parameters, but instead span a range of values throughout their pulsation cycle. For this reason, the validation of our strategy and line list includes non-variable stars with well-established atmospheric parameters. We first tested this line list in a NARVAL